# Mitigating Polarization in Recommender Systems via Network-aware Feedback Optimization

### v1

@@ Abstract
We consider a recommender system that takes into account the interaction between recommendations and the evolution of user interests. Users opinions are influenced by both social interactions and recommended content. We leverage online feedback optimization to design a recommender system that trades-off between maximizing engagement and minimizing polarization. The recommender system is agnostic about users’ opinion, clicking behavior, and social interactions, and solely relies on clicks. We establish optimality and closed-loop stability of the resulting feedback interconnection between the social platform and the recommender system. We numerically validate our algorithm when the user population follows an extended Friedkin–Johnsen model. We observe that network-aware recommendations significantly reduce polarization without compromising user engagement.
@@ 1 Introduction
Online social platforms use recommender systems to provide users with tailored content to maximize engagement over the platform. To do so, recommender systems exploit methods, such as content-based filtering [1] and collaborative filtering [2], that combine information personalization, popularity and similarity of interests with other users. An important drawback of content personalization is related to the amplification phenomenon [3] associated with the feedback loop that arises between recommended content and user preferences. Amplification leads to misinformation spread [4], filter bubbles [5], bias exacerbation [6] and ultimately polarization [7, 8]. Driven by these considerations we propose a recommender system that takes into account the interplay between recommendations and user interests. Our goal is to understand i) the impact of recommendations on user opinions and ii) how should recommendation system depart from engagement maximization to mitigate polarization. By leveraging on online feedback optimization (OFO) [9] and building upon [10, 11], we design a network-aware, multiple objective, recommender system as a feedback controller, illustrated in Figure 1.
@@
Figure 1: Illustration of the closed-loop between the opinion dynamics and the recommender system.
Related Work
Recommendation systems. Most of the works on recommender system are static (e.g. matrix factorization) and neglect the dynamic nature of the feedback interaction between the user and the online platforms. Our work develops along a recent systems-theory perspective on recommender systems [10] where the feedback interaction between social platform and recommender systems is made explicit. In [10], the recommender system provides only extreme positions to the user based on its clicking history. They show that when confirmation bias [12], namely the tendency to engage with content that aligns with prior beliefs, is the only driving force to click on contents, the recommender system induces opinion polarization. The recommender design proposed in [13] alternates between exploitation and exploration by either providing news which more closely align with the users’ beliefs or exploring new content. They show that shifts in the opinions at the individual level do not necessarily reflect in the opinion distribution of the population. The work [14] models the recommender system as an artificial user in the network that provides news content at minimum distance from all users opinions to promote engagement. The authors study how recommender systems of this form affect opinion dynamics. In [11], they consider the interplay between one single user, the recommender and harmful content and seek recommendation policies that tradeoff between maximizing the click-through rate (CTR) and mitigate harm. Their finding suggest that to properly mitigate harm, it is necessary to account for user dynamics. Importantly, most of the works related to recommender systems largely focus on one single user [11, 13, 10] without explicitly considering the influence of social interactions.
Online feedback optimization. Conventional OFO algorithms employ first-order gradient methods [15] that require knowledge about real-time states, sensitivity of the input-output steady-state map and the gradients of the objective function with respect to the input and the states. Recent advancements [16, 17, 18, 19], however allow the employment of OFO with less information. In particular, in [16] the sensitivity matrix is estimated online using online state information. In [17, 18] an OFO scheme is proposed where supervised learning is employed for real-time states and gradient estimation, respectively. Finally, in [19], zeroth-order methods are used to directly estimate the composite gradient using cost function evaluations, thus not requiring sensitivity estimation. See Table 1 for an overview of the recent advancements in OFO.
Table 1: State-of-the-art OFO algorithms.
OFO Method Sensitivity States Cost function Gradients Convexity
[15]
[16]
[17]
[18, Alg. 2]
[19]
Our work
Contributions
Recommender systems provide users tailored content based on implicit feedback, such as clicks. We design an OFO-based recommender system algorithm that extracts information on state, sensitivity and objective function gradient from clicks and dynamically provides user-tailored recommendations.
The recommender system encompasses polarization mitigation and engagement maximization in the objective function. We show that by incorporating information about real-time opinions and clicking behaviors of users as well as the topology of the social network, the recommender is able to enhance engagement while mitigating polarization.
The recommendation algorithm is interpreted as a controller that incorporates opinion (state), topology (sensitivity), and clicking behavior (gradient) estimations. The system dynamics is assumed to be uniformly exponentially stable for constant inputs and to admit a continuously differentiable steady-state map. We provide theoretical closed-loop guarantees between the opinion dynamics and the recommender. Numerical results show that our network-aware recommender system compares favourably in terms of polarization with negligible engagement losses against isolated user scenarios, where the recommendations are solely provided based on personalization and network interactions are neglected.
Our contribution is threefold:
  1. 1.
    We propose a recommender system algorithm that takes into account users’ preferences and trades off between maximizing CTR and minimizing polarization. The recommender system is agnostic about real-time opinions, network interactions and clicking behaviour.
  2. 2.
    From a technical perspective, the OFO algorithm runs combining estimations of system state, input-output sensitivity, and objective function gradient on a non-convex optimization problem. As an independent technical contribution, we provide convergence guarantees towards a locally optimal solution, thereby improving the state of the art on feedback optimization.
  3. 3.
    We validate the proposed control scheme on synthetic data generated via an extension of the Friedkin–Johsnen (FJ) model [10] that also takes into account the recommended content influence over opinions. We show the importance of network awareness on the attenuation of opinion polarization.
Outline
In Section 2, we formulate the optimization problem for engagement maximization and polarization mitigation. In Section 3, we describe the recommender design using OFO. In Section 4, we provide numerical simulations to show our algorithm performances and illustrate the benefits of network-awareness. Section 5 concludes our work.
Notations and Preliminaries
We let the symbols $\mathbb{R}$($\mathbb{R}_{+}$), $\mathbb{N}_{0}$ denote the set of (positive) real numbers and non-negative integers, respectively. The set of integers $\{1,2,\ldots,n\}$ is denoted by $[n]$. For a vector $y\in\mathbb{R}^{n}$, we let the symbol $|y|$ denote the vector whose $i$-th entry, $|y|_{i}$, is the $i$-th component of $y$ in modulus, $|y_{i}|$. The symbol $\langle\cdot,\cdot\rangle:\mathbb{R}^{n}\times\mathbb{R}^{n}\rightarrow\mathbb{R}$ denotes the standard inner product on $\mathbb{R}^{n}$. The value of a signal $x$ at time instant $k$ is denoted by $x^{k}$. The symbol $1_{n},(0_{n})$, denotes the all ones (zeros) vector of size $n$. The symbols $I_{n}$ and $O_{n}$, denote the $n$-dimensional identity and zero matrix, respectively. Given a matrix $A\in\mathbb{R}^{m\times n}$, the matrix is positive (non-negative) if all its elements are greater (or equal) than zero and we denote it as $A>0$ ($A\geq 0$). We let $\mathrm{vec}(A)$ denote the vectorized version of $A$, i.e., if $A=[a_{1}\ a_{2}\ \ldots\ a_{n}]$, with $a_{j}\in\mathbb{R}^{m},j\in[n]$, then $\mathrm{vec}(A)=[a_{1}^{\top}a_{2}^{\top}\ldots a_{n}^{\top}]^{\top}$. For a matrix $A\in\mathbb{R}^{n\times n}$, we denote by $\mathrm{det}[A]$ its determinant; $A$ is row-stochastic if $A\geq 0$ and $A1_{n}=1_{n}$. The symbol $\mathcal{D}_{\mathcal{U}}\subseteq\mathbb{R}$ denotes the set of $n$-dimensional diagonal matrices with diagonal entries in $\mathcal{U}$. The symbol $\mathrm{diag}[x]$ denotes the diagonal matrix with $x_{i}$ on its $i$-th diagonal entry. Given $\mathcal{C}\subset\mathbb{R}^{n}$ and $x\in\mathbb{R}^{n}$, we let $\Pi_{\mathcal{C}}[x]$ denote the Euclidean projection of $x$ over $\mathcal{C}$. The normal distribution is denoted by $\mathcal{N}(\mu,\Sigma)$, with $\mu,\Sigma$ representing the mean and variance, and the uniform distribution over the interval $[a,b]$, with $a,b,\in\mathbb{R}$ is represented by $\mathcal{U}[a,b]$.
@@
The signal $u\in\mathbb{R}^{n}$ is said to be persistently exciting of order $n$ if there exists a time span $S>0$ such that for all $k>0$, the matrix formed by the columns $u^{k+i}$ for $i\in\{0,1,\ldots S\}$ has full rank, i.e. $\mathrm{Rank}[u^{k},u^{k+1},\ldots,u^{k+S}]=n$.
@@
A continuously differentiable map $\Phi:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is globally $\beta$-smooth if its gradient $\nabla\Phi$ is $\beta$-Lipschitz.
@@
Given a globally $\beta$-smooth map $\Phi:\mathbb{R}^{n}\rightarrow\mathbb{R}$, then $\Phi(x_{1})-\Phi(x_{2})-\nabla\Phi^{\top}(x_{2})(x_{1}-x_{2})\leq\frac{1}{2}\beta\lVert x_{1}-x_{2}\rVert^{2}$. Moreover, if $\Phi$ is twice continuously-differentiable, then $\lVert\nabla^{2}\Phi\rVert\leq\beta$.
@@ 2.1 Social Network Modelling
This assumption can be easily adopted for a wide array of renowned opinion dynamics models, such as De-Groot model [23], FJ model [24], and polar opinion dynamics model [25], among others, for which the influence of a recommendation system can be, in principle, included. For an extensive list of such models, we refer the reader to the survey [26]. Traditionally, in fact, these models consider social interactions and prejudice as the only driving forces behind opinion formation. Nonetheless, recommendations on contentious topics provably influence user opinions [10], underscoring the need to include their influence in the opinion dynamics. Next, we illustrate an example of how the FJ model, one of the most widely used opinion dynamics models across different communities [27, 28, 29, 30, 31, 10, 11], can be extended to capture the effects of the recommender on networked users.
@@
Engagement of the users over the platform is measured in terms of CTR, measured as number of clicks over number of provided news articles. As common practice in the context of implicit feedback recommendation systems [32], for a generic user $i\in[n]$ holding an opinion $x_{i}$, we model the probability of $i$ clicking on a recommendation with position $p_{i}$ as a random variable $c_{i}$ sampled from a Bernoulli distribution with unknown argument $g_{i}(p_{i},x_{i})$, i.e. $c_{i}\sim\mathcal{B}(g_{i}(p_{i},x_{i}))$. The argument $g_{i}(p_{i},x_{i})$ represents the clicking behaviour of user $i$ and depends on the user’s opinion $x_{i}$ as well as the position of the recommendation $p_{i}$. We will work under the following regularity assumptions on the clicking behaviour:
@@
The clicking behaviour $g(p,x):[-1,1]^{n}\times[-1,1]^{n}\to[0,1]^{n}$ is $M_{x}$-Lipschitz with respect to $x$, and $L_{p}$- and $L_{x}$-smooth with respect to $p$ and $x$, respectively. $\square$
A concrete example of a clicking behaviour is provided.
@@
A user $i\in[n]$, holding opinion $x_{i}$, affected by extremity confirmation bias clicks on a recommendation $p_{i}$ with probability
@@
where the clicking behaviour $g_{i}(p_{i},x_{i})=\frac{1}{2}+\frac{1}{2}x_{i}p_{i}$ models confirmation bias towards extreme recommendations. In fact, if the opinion $x_{i}\approx\pm 1$ and the position $p_{i}\approx\pm 1$ ($\mp 1$), then the probability of clicking is almost $1$ (00). The click event becomes random as the user position approaches the neutral stance $p_{i}\approx 0$, highlighting diminished engagement for less polarized content or when recommendations directly counter the user’s stance on the issue. $\triangle$
@@ 2.2 Problem formulation
The goal of the recommender system is to provide recommendations that optimize a specific metric, denoted as $\varphi(p,x)$. Typical choices include maximizing CTR [10, 11], opinion control [33], and polarization mitigation [34]. Deviating from traditional recommendation systems, we consider a multi-objective cost function combining engagement maximization and polarization mitigation, given by
@@
where $\gamma$ is a non-negative trade-off parameter. The engagement-related term in the cost (3) is the CTR, formalized as
@@
where $\mathbb{E}_{c_{i}\sim\mathbb{B}(g_{i}(x_{i},p_{i}))}[\,c_{i}\,]$ is the expectation of user $i$’s clicking, given their opinion $x_{i}$ and a recommendation with position $p_{i}$. The second term in (3), $\varphi^{\rm pol}$, is a smooth Lipschitz function that measures opinion polarization. For our algorithm we consider it to be $\varphi^{\text{pol}}(x):=\sum_{i\in[n]}s_{i}(x)$, where $s_{i}$ is a soft penalty function:
@@
Therein, the parameters $\epsilon_{1}\leq\epsilon_{2}$ are used to control the degree of penalty towards extreme opinions. Specifically, a smaller positive (negative) choice for $\epsilon_{2}$ ($\epsilon_{1}$) indicates a higher penalty on extreme positive (negative) opinions. According to [35], the metric in (4) falls in the class of polarization spread, indicating how extreme the opinions are. Note that, for $\epsilon_{1}=\epsilon_{2}=0$, the term (4) reduces to $\varphi^{\text{pol}}(x)=\lVert x\rVert^{2}$, a classic measure for opinion polarization [28].
@@
This problem is uncertain and potentially non-convex, since the analytical form of the clicking behaviour is unknown and likely complex. Likewise, the opinion dynamics (1) are unknown. If the opinion dynamics in (1) and the clicking behavior $g_{i}$ were known, and an accurate prediction of external influences $d$ were available, then one could aim to solve (5) offline. However, in practice, none of these information is readily available, and the recommender system provides positions solely based on clicks $\{c_{i}^{k}\}$.
@@
Equation (5b) translates into the users’ opinion stationarity before the recommender system action is taken. This is consistent with most of the works on recommender systems based on preference dynamics[11, 10, 36, 6, 37], where the action is taken based on the steady-state dynamics of the proposed model.
@@
Design a feedback controller $p(c)$ so that (1) converges to a solution $(p^{*},x^{*})$ of the optimization problem (5), assuming that only clicks $c_{i}^{k}$ are available. The clicking behaviour $g_{i}(p_{i},x_{i})$, steady-state map $h(p,d)$, the opinion dynamics (1), external influence $d$ are unknown, and the real-time opinions $x$ are not measurable.

### v2

# Network-aware Recommender System via Online Feedback Optimization

@@ Abstract
Personalized content on social platforms can exacerbate negative phenomena such as polarization, partly due to the feedback interactions between recommendations and the users. In this paper, we present a control-theoretic recommender system that explicitly accounts for this feedback loop to mitigate polarization. Our approach extends online feedback optimization – a control paradigm for steady-state optimization of dynamical systems – to develop a recommender system that trades off users engagement and polarization reduction, while relying solely on online click data. We establish theoretical guarantees for optimality and stability of the proposed design and validate its effectiveness via numerical experiments with a user population governed by Friedkin–Johnsen dynamics. Our results show these “network-aware” recommendations can significantly reduce polarization while maintaining high levels of user engagement.
@@ 1 Introduction
Online social platforms use recommender systems to provide users with tailored content to maximize engagement over the platform. To do so, recommender systems exploit techniques, such as content-based filtering [1] and collaborative filtering [2], that combine information personalization, popularity and similarity of interests with other users. A significant drawback of content personalization is related to the amplification phenomenon [3], associated with the feedback loop that arises between recommended content and user preferences. Amplification leads to misinformation spread [4], filter bubbles [5], bias exacerbation [6] and, ultimately, polarization [7, 8]. Motivated by these challenges, we consider instead recommender systems inspired by control-theoretic principles that have the potential of explicitly accounting for the feedback interplay between recommendations and user interests to mitigate polarization. Our goals are to develop a systems-theory-based recommender design to understand i) the impact of recommendations on user opinions, and ii) how recommender systems should depart from engagement maximization to mitigate polarization. We approach the problem by building upon the systems-theory perspective of recommender systems pioneered in [9, 10] and by leveraging on online feedback optimization (OFO) [11] – an emerging control paradigm for steady-state optimization of dynamical systems. Specifically, we aim at designing a recommender system as a dynamic feedback controller that mitigates polarization by providing users personalized content, using solely implicit feedback from their online click data, as illustrated in Figure 1.
Figure 1: Closed-loop interconnection between users’ opinion in a social network and the proposed recommender system.
Related Work
Recommender systems. Most of the literature on recommender systems relies on static methods (such as matrix factorization) and neglects the dynamic nature of the feedback interactions between the users and the online platforms [12]. In contrast, our work develops along a more recent systems-theory perspective on recommender systems [9, 10], where the feedback interaction between social platform and recommender systems is made explicit. In [9], the recommender system provides only extreme positions to the users based on their clicking history. The authors show that when confirmation bias [13] – the tendency to engage with content that aligns with prior beliefs – is the sole factor driving clicks, their recommender system promotes opinion polarization. The recommender design proposed in [14] alternates between exploitation and exploration by either providing news which more closely align with the users’ beliefs or exploring new content. The authors show that shifts in the opinions at the individual level do not necessarily reflect in the opinion distribution of the population. In [15], the recommender system is modeled as an artificial user within the social platform, providing news content at minimum distance from all users opinions to promote engagement. This study investigates how such a recommender system can affect opinion dynamics across the network. In [10], the focus is on the interplay between a single user, the recommender system, and harmful content. The authors seek recommendation policies that balance maximizing the click-through rate (CTR) and mitigate harm. Their findings suggest that to properly mitigate harm, it is necessary to account for user dynamics. Notably, most of these studies focus solely on individual-user designs, where recommendations are tailored to each user in isolation, without explicitly accounting for the broader influence of social interactions. However, social interactions are well-known to be a crucial factor in shaping opinions over a connected population [16, 17, 18].
Online feedback optimization. Most OFO controllers are based on first-order optimization algorithms [11], and their deployment require real-time access to the states of the plant, an explicit knowledge of its input-output sensitivity (or a closed approximation), and a closed-form expression of the gradient of the control objective. However, some recent extensions [19, 20, 21, 22, 23] have relaxed some of these stringent requirements. Specifically, [19] proposes to estimate the input-output sensitivity of the plant from real-time state measurements using recursive-least square; while [20] achieves similar results leveraging historical input-output trajectories and behavioral systems theory. In [21], the authors address the case where exact states measurements cannot be directly accessed, but need to be inferred from perceptual information. Their solution relies on the integration of a neural network component that transform high-dimensional sensory data into state estimates. Similarly, [22] proposes a neural-network solution for the case where a closed-form expression of the gradient is not available. Finally, [23] utilizes a zeroth-order method to estimate the composite gradient directly from online cost function evaluations, eliminating the need for both input-output sensitivity and a closed-form gradient expression. See Table 1 for a comparative overview of these OFO controllers.
Table 1: Comparative overview of the main requirements of some recent online feedback optimization (OFO) controllers.
Controller Sensitivity States Objective Gradient Convex
[11]
[19, 20]
[21]
[22, Alg. 2]
[23]
This work
Contributions
Unfortunately, existing OFO designs cannot be directly applied to social networks, as recommender systems typically only have access to implicit feedback in the form of click data. Real-time users opinion are inaccessible, social interactions may be hidden due to privacy preferences, and there is no closed-form expression for the expected CTR as it depends on unknown users clicking behavior – namely, their tendencies or pattern when interacting with recommended items.
In this paper, we address these challenges by integrating auxiliary estimation blocks into a projected-gradient OFO controller to compensate for the missing information, thereby enabling online feedback optimization in social networks. Specifically, we estimate opinions and clicking behaviour using supervised learning, sensitivity using Kalman filtering and gradients using zero-order methods.
Unlike most system-theory-based recommender system designs in the literature, our proposed algorithm is network-aware, namely, it leverages on users’ social interactions to better mitigate undesired population phenomena such as polarization, while maintaining high levels of engagement.
Our contribution is threefold:
  1. (i)
    We propose a novel OFO-based recommender system algorithm that steers a social network towards steady-state opinion profiles that balance between maximizing overall click-through rate and minimizing polarization. Our design is agnostic about real-time opinions of the users, their network interactions and clicking behavior, and solely relies on click data measured online.
  2. (ii)
    We establish closed-loop stability guarantees of the recommender system-social network interconnection with respect to locally-optimal solutions of the steady-state control objective, assuming exponential stability of the opinion dynamics and other mild technical conditions.
  3. (iii)
    We validate the proposed recommender system through numerical simulations on a social network where user opinions are governed by an extended version of the Friedkin–Johnsen (FJ) dynamics [9], which incorporates the influence of recommended content. Our results demonstrate that network-awareness plays a crucial role in reducing opinion polarization in social platforms.
Outline
In Section 2, we formulate the recommender objectives as a steady-state control problem. In Section 3, we describe the recommender design using OFO. Section 4 present numerical simulations to evaluate the performance of our algorithm and highlight the benefits of network-awareness. Finally, Section 5 summarizes our findings and concludes the paper.
Notations and Preliminaries
@@
The signal $u\in\mathbb{R}^{n}$ is said to be persistently exciting of order $n$ if there exists a time span $S>0$ such that for all $k>0$, the matrix formed by the columns $u^{k+i}\in\mathbb{R}^{n}$ for $i\in\{0,1,\ldots S\}$ has full rank, i.e. $\mathrm{Rank}[u^{k}\ u^{k+1}\ \ldots\ u^{k+S}]=n$.
@@
A continuously differentiable map $\Phi:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is globally $\beta$-smooth if its gradient $\nabla\Phi$ is globally $\beta$-Lipschitz.
@@
Given a globally $\beta$-smooth map $\Phi:\mathbb{R}^{n}\rightarrow\mathbb{R}$, then $\Phi(x_{1})-\Phi(x_{2})-\nabla\Phi^{\top}(x_{2})(x_{1}-x_{2})\leq\frac{1}{2}\beta\lVert x_{1}-x_{2}\rVert^{2}$. Further, if $\Phi$ is twice continuously-differentiable, then $\lVert\nabla^{2}\Phi\rVert\leq\beta$.
@@ 2.1 Social Network Modelling
This assumption can be easily adopted for a wide array of renowned opinion dynamics models, such as De-Groot model [17], FJ model [16], and polar opinion dynamics model [18], among others, for which the influence of a recommender system can be, in principle, included. For an extensive list of such models, we refer the reader to the survey [27]. Traditionally, in fact, these models consider social interactions and prejudice as the only driving forces behind opinion formation. Nonetheless, recommendations on contentious topics provably influence user opinions [9], underscoring the need to include their influence in the opinion dynamics. Next, we illustrate an example of how the FJ model, one of the most widely used opinion dynamics models across different communities [28, 29, 30, 31, 32, 9, 10], can be extended to capture the effects of the recommender on networked users.
@@
Engagement of the users over the platform is measured in terms of CTR, measured as number of clicks over number of provided news articles. As common practice in the context of implicit feedback recommender systems [33], for a generic user $i\in[n]$ holding an opinion $x_{i}$, we model the probability of $i$ clicking on a recommendation with position $p_{i}$ as a random variable $c_{i}$ sampled from a Bernoulli distribution with unknown argument $g_{i}(p_{i},x_{i})$, i.e. $c_{i}\sim\mathcal{B}(g_{i}(p_{i},x_{i}))$. The argument $g_{i}(p_{i},x_{i})$ represents the clicking behaviour of user $i$ and depends on the user’s opinion $x_{i}$ as well as the position of the recommendation $p_{i}$. We will work under the following regularity assumptions on the clicking behaviour:
@@
The clicking behaviour $g(p,x):[-1,1]^{n}\times[-1,1]^{n}\to[0,1]^{n}$ is $M_{x}$ Lipschitz with respect to $x$, and $L_{p}$- and $L_{x}$-smooth with respect to $p$ and $x$, respectively. $\square$
Next, we provide an example of a clicking behaviour.
@@
A user $i~{}\in~{}[n]$, holding opinion $x_{i}$, affected by extremity confirmation bias clicks on a recommendation $p_{i}$ with probability
@@
where the clicking behaviour $g_{i}(p_{i},x_{i})=\frac{1}{2}+\frac{1}{2}x_{i}p_{i}$ models confirmation bias towards extreme recommendations. In fact, for an opinion $x_{i}\approx\pm 1$ and a position $p_{i}\approx\pm 1$ ($\mp 1$), the probability of clicking is almost $1$ (00). The click event becomes random as the user position approaches the neutral stance $p_{i}\approx 0$, highlighting diminished engagement for less polarized content or when recommendations directly counter the user’s stance on the issue. $\triangle$
@@ 2.2 Problem formulation
The goal of the recommender system is to provide recommendations that optimize a specific metric, denoted as $\varphi(p,x)$. Typical choices in the literature include maximizing CTR [9, 10], opinion control [34], and polarization mitigation [35]. Deviating from traditional recommender systems, we consider a multi-objective cost function combining engagement maximization and polarization mitigation, given by
@@
where $\gamma$ is a non-negative trade-off parameter. The engagement-related term in (3) is the CTR, formalized as
@@
where $\mathbb{E}_{c_{i}\sim\mathbb{B}(g_{i}(x_{i},p_{i}))}[\,c_{i}\,]$ is the expectation of user $i$’s clicking, given their opinion $x_{i}$ and a recommendation with position $p_{i}$. The second term in (3), $\varphi^{\rm pol}$, is a smooth and Lipschitz function that measures opinion polarization. For our algorithm we consider it to be $\varphi^{\text{pol}}(x):=\sum_{i\in[n]}s_{i}(x)$, where $s_{i}$ penalizes deviations from the moderate (zero) opinions:
@@
Therein, the parameters $\epsilon_{1}\leq\epsilon_{2}$ are used to control the degree of penalty towards extreme opinions. Specifically, a smaller positive (negative) choice for $\epsilon_{2}$ ($\epsilon_{1}$) indicates a higher penalty on extreme positive (negative) opinions. According to [36], the metric in (4) falls in the class of polarization spread, indicating how extreme the opinions are. Note that, for $\epsilon_{1}=\epsilon_{2}=0$, the term (4) reduces to $\varphi^{\text{pol}}(x)=\lVert x\rVert^{2}$, another classic measure for opinion polarization [29].
@@
This problem is uncertain and potentially non-convex, since the analytical form of the clicking behaviour is unknown and likely complex. Likewise, the opinion dynamics (1) are unknown. If the opinion dynamics in (1) and the clicking behavior $g_{i}$ were known, and an accurate prediction of external influences $d$ were available, then one could aim to solve (5) offline. However, in practice, none of these information is readily available, and the recommender system can provide personalized positions solely relying on clicks $\{c_{i}^{k}\}$.
@@
Equation (5b) reflects the assumption that users’ opinions reach steady-state before the recommender system takes action. This approach is in line with most of the existing works on recommender systems that incorporates preference dynamics [10, 9, 37, 6, 38], where recommendation decisions are made based on the steady-state behavior of the users’ preferences under the model.
@@
Design a feedback controller $p(c)$ so that (1) converges to a solution $(p^{*},x^{*})$ of the optimization problem (5), assuming that only clicks $c_{i}^{k}$ are available. The clicking behaviours $g_{i}(p_{i},x_{i})$, steady-state map $h(p,d)$, the opinion dynamics (1), external influence $d$ are unknown, and the real-time opinions $x$ are not measurable.